# Object-centric proto-symbolic behavioural reasoning from pixels

### v2

@@ 7.1 Planning
${\bm{\mathbf{\pi}}}^{*}=({\bm{\mathbf{U}}}^{T}{\bm{\mathbf{L}}}{\bm{\mathbf{U}}}+\lambda_{a}{\bm{\mathbf{I}}})^{-1}{\bm{\mathbf{U}}}^{T}{\bm{\mathbf{L}}}\left({\bm{\mathbf{1}}}\otimes(\tilde{{\bm{\mathbf{\mu}}}}-{\bm{\mathbf{\mu}}}_{s^{\dagger}_{t}})-{\bm{\mathbf{d}}}^{(T)}\otimes\begin{bmatrix}{\bm{\mathbf{\mu}}}_{s^{\prime}_{t}}\\ 0\end{bmatrix}\right)$ (21)
@@ 8 Latent space traversals
OBR is trained without supervision, and so the structure of its latent representational space is a priori unknown. Here, we explore this space by systematically traversing it one dimension at a time. We first let OBR encode an image into its latent space. Subsequently, for each object representation, we increment the target dimension’s encoded (mean) value with a range of 20 evenly spaced values between $[-1,1]$. We then feed the resulting, perturbed latent vectors through OBR’s decoder to obtain 20 images that vary systematically (and for all objects) along the target dimension. The images in Fig. 10 show the results of this procedure for three different scenes, for each of the 16 latent dimensions of the trained OBR model. Of particular note are dimensions 3 and 9, which have learned to encode the objects’ positions (along an approximately orthogonal set of axes at an arbitrary angle to the pixel grid). Other recognizable variations can be seen in color (dimensions 4, 6 and 11) and size (14). Shape and orientation appear to be entangled along multiple dimensions. Other dimensions do not obviously encode anything – they may be truly non-coding, or their role might not be visible in combination with the other latent values that we happened to sample here. In particular, the depth value of an object is almost certainly encoded in one or more latent dimensions, but this is not apparent from these traversals, as they target all objects simultaneously, and thus should not affect their occlusion patterns (which object is in front of which other object(s)).

### v3

@@ 7.1 Planning
[…]
@@
7.1.1 Baseline configurations
All ablation experiments and baseline comparisons were conducted on a single NVIDIA H200 (141 GB) with PyTorch 2.7 and CUDA 12.8. SAVi variants were trained from scratch using Adam optimizer (lr = $1\times 10^{-4}$ with cosine decay to $1\times 10^{-6}$) on 50k Active-dSprites sequences with 4 slots (128-dim). V-JEPA variants share a frozen ViT-L encoder (1024-dim), 12-layer predictor (512-dim), and action head (12-dim). Frame stacking (1–8 frames) and learning rates ($10^{-3}$$10^{-6}$) were systematically explored. Optimal configurations: V-JEPA Original (4 frames, L1 + rollout loss), V-JEPA BC (1 frame, MSE loss), V-JEPA DAgger (2 frames, MSE loss with iterative data aggregation, $\beta=0.8$ decay). Dynamics module ablations (linear, GRU, GateL0RD, LSTM, Transformer) were trained on 8k state trajectories with batch sizes optimized per architecture.
Training duration. SAVi variants: 50 epochs. 1-frame variant approximately 1 hour; 4-frame variant required approximately 3.5 hours, 8-frame variant approximately 6.5 hours; and 16-frame variant required approximately 12 hours. V-JEPA 2-AC Original: 100 epochs (6 hours), did not converge. V-JEPA 2-AC BC: 50 epochs (3.2 hours), converged at epoch 26 (1.5 hours). V-JEPA 2-AC DAgger: 50 epochs across 6 data aggregation iterations (6 hours total), converged at epoch 22 (2 hours). Dynamics modules: 50 epochs (1–3 hours per architecture depending on model complexity).
@@ 8 Latent space traversals
OBR is trained without supervision, and so the structure of its latent representational space is a priori unknown. Here, we explore this space by systematically traversing it one dimension at a time. We first let OBR encode an image into its latent space. Subsequently, for each object representation, we increment the target dimension’s encoded (mean) value with a range of 20 evenly spaced values between [1,1][-1,1]. We then feed the resulting, perturbed latent vectors through OBR’s decoder to obtain 20 images that vary systematically (and for all objects) along the target dimension. The images in Fig. 15 show the results of this procedure for three different scenes, for each of the 16 latent dimensions of the trained OBR model. Of particular note are dimensions 3 and 9, which have learned to encode the objects’ positions (along an approximately orthogonal set of axes at an arbitrary angle to the pixel grid). Other recognizable variations can be seen in color (dimensions 4, 6 and 11) and size (14). Shape and orientation appear to be entangled along multiple dimensions. Other dimensions do not obviously encode anything – they may be truly non-coding, or their role might not be visible in combination with the other latent values that we happened to sample here. In particular, the depth value of an object is almost certainly encoded in one or more latent dimensions, but this is not apparent from these traversals, as they target all objects simultaneously, and thus should not affect their occlusion patterns (which object is in front of which other object(s)).